OpenStudy (anonymous):

Quadrilateral PAID is a rectangle whose diagonals have the endpoints P(-3, -2)I(4, -7) and A(4, -2)D(-3, -7). Find the diagonals' intersection point. In complete sentences, explain which method you used for finding the intersection point. Complete your work in the space provided or upload a file that can display math symbols if your work requires it. If necessary, leave your answers in terms of a fraction.

OpenStudy (anonymous):

I need help on this

OpenStudy (anonymous):

it tell me to write a complete sentences

OpenStudy (owlcoffee):

Well, you would have to find the equation of the line that represents the diagonals, and then, just intersect them with a system of equation 2x2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@rational

OpenStudy (rational):

you may simply find the midpoint of PI

OpenStudy (anonymous):

I did but tell me to write a complete sentence about this

OpenStudy (rational):

P=(-3, -2) I=(4, -7) Since the diagonals bisect each other in a rectangle, the diagonals intersect at the midpoint of diagonals. use the midpoint formula to find the intersecting point of diagonals

OpenStudy (rational):

midpoint = \[\left(\frac{-3+4}{2},~~\frac{-2-7}{2}\right)=\left(\frac{1}{2},~~\frac{-9}{2}\right)\]

OpenStudy (anonymous):

ok I solve all

OpenStudy (rational):

you're done! midpoint is same as the intersecting point of diagonals

OpenStudy (anonymous):

yes